OpenStudy (ttp):

For breakfast, the cafeteria offers scrambled eggs with one slice of toast or an egg sandwich made with two slices of toast. In fifteen minutes, they sold 28 breakfasts and used 47 slices of bread. Use a model to determine how many egg sandwiches they sold. 16 17 18 19

OpenStudy (lexipoo1998):

16

OpenStudy (lexipoo1998):

16 * 2 = 32 47-32 = 15 15+32 = 47 So it has be 16